OpenStudy (anonymous):

how to solve x^2 + 4x +3

OpenStudy (anonymous):

solve as in "set it equal to zero and solve"?

OpenStudy (anonymous):

yea! ;o

OpenStudy (anonymous):

(-b+_(b^2-4ac)^1/2)/2a

OpenStudy (anonymous):

can u do that with numbers please lol sorry ..all dat jus made my brain crash

OpenStudy (anonymous):

If you set this expression equal to 0, you have equation x^2 + 4x + 3 = 0 which is in the form ax^2 + bx + c = 0, so your a=1, b=4, and c=3. Rewriting the equation from another poster:\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]will be your solution(s). Just substitute.